OpenStudy (anonymous):

a • b • 4 = 4 • a • b is an example of which algebraic property? Associative Property of Multiplication Transitive Property Reflexive Property Commutative Property of Multiplication

OpenStudy (anonymous):

Commutative Property of Multiplication

OpenStudy (anonymous):

ummm... Associative Property of Multiplication

OpenStudy (anonymous):

no .. associative means a*(b*c) = (a*b)*c .. its not associative its commutative.

OpenStudy (anonymous):

no, commutative deals with 2 numbers while associative deals with three or more

OpenStudy (anonymous):

let a*b = u then 4 * u = u*4 4 * a * b = a * b * 4

OpenStudy (anonymous):

there you go .. two numbers same question

OpenStudy (anonymous):

*face palm*